OpenStudy (anonymous):

Can someone help me to find a vertex of parabola??? y=4x^2 + 4x -5

OpenStudy (xapproachesinfinity):

\[\text{the technique to find to vertex of the probola given it's standard equation }\\ \text{y=ax^2+bx+c is called completing the square}\]

OpenStudy (anonymous):

not the way i do it

OpenStudy (xapproachesinfinity):

so you need to complete the square for y=4x^2+4x-5

OpenStudy (xapproachesinfinity):

@satellite73 you mean u use calculus?

OpenStudy (xapproachesinfinity):

@Savannah_Lynn13 hey there interested to do this?

OpenStudy (misty1212):

HI!!

Nnesha (nnesha):

no he use `-b/2a` formula :D

OpenStudy (misty1212):

first coordinate of the vertex is always \(-\frac{b}{2a}\) and if you know the first coordinate you find the second by substitution

Nnesha (nnesha):

no he use `-b/2a` formula :D to find x-coordinate of the vertex and then substitute all x's for its value into the equation to find y-coordinate :P

OpenStudy (misty1212):

here we have \(a=4,b=4\) put in in to \[\huge-\frac{b}{2a}\] to get the \(x\) value of the vertex

OpenStudy (xapproachesinfinity):

well that's if he he knows that (-b/2a, f(-b/2a) ) is the vertex

OpenStudy (anonymous):

ok so the thing i plug in would be -1/2, correct?

OpenStudy (misty1212):

yes

OpenStudy (xapproachesinfinity):

yes x=-1/2 you need f(-1/2) too

OpenStudy (xapproachesinfinity):

i prefer the other way which show where thing come from it is okay to use x=-b/2a as long as you know where it comes from

OpenStudy (xapproachesinfinity):

i though @satellite73 meant to take first derivative lol when he not the way i do it

OpenStudy (shadowlegendx):

I think I skipped class that day

OpenStudy (xapproachesinfinity):

hahah lol

OpenStudy (xapproachesinfinity):

still need help i guess you know thee answer

OpenStudy (shadowlegendx):

oh completing the square

OpenStudy (anonymous):

so i get to\[y=4(-\frac{ 1 }{2})^2+4(-\frac{ 1 }{ 2 })-5\] and no i don't know the anwer

OpenStudy (anonymous):

it is arithmetic from here on in

OpenStudy (anonymous):

what is \(\left(-\frac{1}{2}\right)^2\)?

OpenStudy (anonymous):

yeah and i have a very difficult time with math, so for some reason this is really confusing me

OpenStudy (anonymous):

if you don't know, i will tell you, but it is easy to square a fraction, square the top and the bottom

OpenStudy (anonymous):

-1/4

OpenStudy (anonymous):

no, when you square the minus sign goes away

OpenStudy (anonymous):

\[\left(-\frac{1}{2}\right)^2=\left(-\frac{1}{2}\right)\times\left(-\frac{1}{2}\right)=\frac{1}{4}\]

OpenStudy (anonymous):

now we are at \[y=4(\frac{ 1 }{4})+4(-\frac{ 1 }{ 2 })-5\]

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

is it -6?

OpenStudy (anonymous):

Is this correct?

OpenStudy (lochana):

y=4x^2 + 4x -5 y = 4(x^2 + x) -5 y = 4(x+1/2)^2 -4(1/4) -5 y = 4(x + 1/2)^2 - 6 vertex point = (-1/2, -6)

OpenStudy (anonymous):

Thank you :)

OpenStudy (lochana):

you need to convert the equation into vertex form as mentioned in that web page. hope it helps